how do you do... substitution on this problem 4p+2q=8
q=2p+1

Answer by KnightOwlTutor(293)

4p+2q=8
Substitute q=2p+1 in the equation above
4p+2(2p+1)=8
Use the distributive property
4p+4p+2=8
8p+2=8
Subtract 2 from both sides
8p=6
divide both sides by 8
p=6/8 or 3/4
Plug p in to get the value for q
q=2(3/4)+1=3/2+2/2=5/2
Plug in p and q values into the equation
4(3/4)+2(5/2)=8
After cancelling terms you get
3+5=8